A company estimates that 0.5% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $200.
If they offer a 2 year extended warranty for $27, what is the company's expected value of each warranty sold?

Answer by ikleyn(52748)

.
A company estimates that 0.5% of their products will fail after the original warranty period
but within 2 years of the purchase, with a replacement cost of $200.
If they offer a 2 year extended warranty for $27, what is the company's expected value
of each warranty sold?
~~~~~~~~~~~~~~~~

Notice that 0.5% = 0.005.


The expected profit of selling each warranty is

    27 - 0.005*200  dollars = 27 - 1 = 26  dollars.


Here $27 is what the customer pays for the 2-years warranty,
and 0.005*200 = 1 dollar is the company's average expense per unit to cover the replacement.


The net  27 - 1 = 26 dollars is the expected company's profit of each warranty sold.


$26 is the price which the customer does agree to pay to the company for peace in mind.

Solved and explained.

----------------

To me,  the presented input data seems to be strange, unfair and unrealistic.

Double check if  0.5%  is correct value in your post.
It is more likely that it is  5%.